# Holmes: Towards Effective and Harmless Model Ownership Verification to Personalized Large Vision Models via Decoupling Common Features

### v1

@@ Abstract
Large vision models achieve remarkable performance in various downstream tasks, primarily by personalizing pre-trained models through fine-tuning with private and valuable local data, which makes the personalized model a valuable intellectual property for its owner. Similar to the era of traditional DNNs, model stealing attacks also pose significant risks to these personalized models. However, in this paper, we reveal that most existing defense methods (developed for traditional DNNs), typically designed for models trained from scratch, either introduce additional security risks, are prone to misjudgment, or are even ineffective for fine-tuned models. To alleviate these problems, this paper proposes a harmless model ownership verification method for personalized models by decoupling similar common features. In general, our method consists of three main stages. In the first stage, we create shadow models that retain common features of the victim model while disrupting dataset-specific features. We represent the dataset-specific features of the victim model by the output differences between the shadow and victim models. After that, a meta-classifier is trained to identify stolen models by determining whether suspicious models contain the dataset-specific features of the victim. In the third stage, we conduct model ownership verification by hypothesis test to mitigate randomness and enhance robustness. Extensive experiments on benchmark datasets verify the effectiveness of the proposed method in detecting different types of model stealing simultaneously.
Index Terms:
Ownership Verification, Model Watermarking, Model Stealing, Deep IP Protection, AI Security
I Introduction
@@
Pre-trained large vision models (LVMs) are widely used across various applications [1, 2, 3, 4, 5, 6]. The dominant development paradigm for specific downstream tasks involves personalizing LVMs via fine-tuning on local data. This approach results in a well-performing task-oriented model at a small training cost since the foundation model has already learned a high-quality representation from a large amount of data [7, 8, 5, 6]. Generally, these personalized models have significant commercial value, and their fine-tuning process involves private and valuable data, making them critical intellectual property for the owners.
@@
However, recent studies revealed that the adversary could ‘steal’ (personalized) models even without access to their training samples or model parameters [9, 10, 11, 12, 13] via model stealing attacks. For example, the adversary might use the victim model to label an unlabeled dataset or apply knowledge distillation to reproduce the model [11, 13]. Such attacks pose a severe threat to model owners, as they require significantly fewer resources than directly conducting personalization ($e.g.$, fewer computational resources or less training data).
@@
Currently, model ownership verification (MOV) is an important audit strategy to protect the model copyright. It aims to identify whether a suspicious third-party model is stolen from the victim. Existing MOV methods usually fall into two categories: model watermarking and model fingerprinting. Specifically, the model watermarking approaches [14, 15, 16, 17, 18, 19] mostly rely on embedding defender-specified misclassification behaviors, especially via backdoor attacks [20, 21, 22, 23], into the victim model. However, these methods inevitably introduce potential vulnerabilities, compromising model reliability and limiting practical deployment. Most recently, to address this limitation, MOVE [24] proposed a harmless watermarking scheme that avoids embedding misclassification behaviors. Specifically, MOVE defined harmlessness as the watermarked model should have similar prediction behaviors to the one without any watermark ($i.e.$, benign model trained with the benign dataset). Accordingly, instead of leveraging misclassification behaviors, MOVE modified a subset of training samples using style transfer without manipulating their labels to embed external features as dataset-specific features into the training set. By identifying whether the model has learned the external features, MOVE can provide effective model ownership verification. In particular, since the modifications to the training set do not alter the labels of the affected data, this method remains harmless. However, it is primarily designed for models trained from scratch. As we will reveal in Section III, MOVE is ineffective for fine-tuned models because the required external features cannot be adequately learned during fine-tuning due to interference from extensive ‘robust features’ learned by the foundation model.
@@
In contrast, model fingerprinting methods [25, 26, 27] verify ownership by detecting whether the suspicious models retain the inherent features naturally learned by the victim model from its training data. For example, dataset inference [25] uses the decision boundary of the victim model as a representation of its learned features. Unlike watermarking, fingerprinting does not modify the model’s training process, therefore naturally ensuring harmlessness. However, these methods are primarily designed for simple model stealing attacks ($e.g.$, direct copying or fine-tuning), whereas exhibiting limited robustness against advanced stealing methods [25, 24]. Besides, as we will show in Section III, relying solely on inherent features may lead to misjudgment when the victim and suspicious models are trained on similar datasets, as they may learn similar ‘common features’. These limitations raise an intriguing question: Can we design effective and harmless MOV methods for personalized LVMs?
@@
Fortunately, the answer to the above question is positive, although its solution is non-trivial. Motivated by the above findings and understandings, this paper proposes to exploit inherent yet dataset-specific features for verification. Intuitively, inherent features can be effectively learned, while the dataset-specific nature reduces the risk of misjudgment. In general, our method (dubbed ‘Holmes’) consists of three main stages, including (1) creating shadow models, (2) training ownership meta-classifier, and (3) ownership verification with hypothesis test. Specifically, in the first stage, we create two shadow models, including one poisoned shadow model and one benign shadow model, to decouple the common features and distinguish the learned dataset-specific features of the victim model. In general, the poisoned shadow model retains common features learned by the victim model, while its dataset-specific features are disrupted. As such, we can indirectly capture the dataset-specific features while decoupling the common features via the output difference between these two models. To achieve it, we poison the most well-learned personalization samples (with the lowest loss value) to disrupt dataset-specific features via the label-inconsistent backdoor attack ($e.g.$, BadNets [28]). Intuitively, we choose these samples since they correspond to the model’s propensity to memorize dataset-specific patterns (instead of common features) [29]. We exploit the (label-inconsistent) backdoor attack since it allows a high performance of poisoned models in predicting benign testing samples, thus preserving common features. The benign shadow model is introduced for reference to further mitigate false positives from independent models with related dataset-specific features. In general, this model exhibits dataset-specific features that are similar but distinct from those of the victim model. Accordingly, the output difference between the benign and poisoned shadow models reveals dataset-specific features that are distinct from those of the victim model. To achieve this model, we fine-tune the same foundation model on personalization samples excluding those that were well-learned by the victim model. In the second stage, we train a binary ownership meta-classifier for identifying dataset-specific features of the victim model, based on two output differences: (1) the output differences between the victim model and the poisoned shadow model (for capturing the victim model’s dataset-specific features) and (2) the output differences between the benign shadow model and the poisoned shadow model (for capturing features unrelated to the victim model’s dataset-specific features). In the third stage, we conduct model ownership verification via hypothesis test to mitigate randomness and further enhance robustness.
@@
The main contribution of this work is three-fold: (1) We revisit the ownership verification methods and reveal the limitations of existing methods for personalized LVMs. (2) We propose an effective and harmless ownership verification method via decoupling common features. (3) We verify the effectiveness of our method on benchmark datasets in detecting various types of model stealing simultaneously.
II Background and Related Work
II-A Large Vision Model
@@
By learning high-quality representations and features from extensive datasets, large pre-trained vision models (LVMs) have demonstrated remarkable capabilities in comprehending image content and extracting rich semantic information [1, 2, 3, 4, 5, 6]. These models excel across a wide range of computer vision tasks, including image classification [30, 31, 32, 33], transfer learning [7, 8, 5, 6], object detection [34, 35, 36], and object segmentation [37, 38, 39]. Among the most notable LVMs, CLIP (Contrastive Language-Image Pre-Training) [2] stands out as a cornerstone in the field. CLIP leverages contrastive learning to align textual and visual information in a shared embedding space, enabling it to perform image classification [30, 31, 32, 33] and transfer learning [40] with exceptional accuracy. Pre-trained on a massive dataset of image-text pairs, CLIP is widely adapted as a foundation model for various downstream tasks [41, 42, 43]. Other prominent LVMs include BEiT (Bidirectional Encoder Representation from Image Transformers) [3], which also achieves outstanding performance in image classification and transfer learning [5].
@@
Currently, the most prevalent application paradigm for LVMs involves personalization on downstream tasks via fine-tuning [4, 7, 8, 5, 6]. This process typically begins with a pre-trained foundation model, which has already learned high-quality generalizable features from a large dataset. The foundation model is then adapted to specific tasks by fine-tuning on task-specific datasets. For example, in medical imaging, a pre-trained LVM can be fine-tuned on a dataset of X-ray images to significantly improve disease diagnosis accuracy [42, 43]. Fine-tuning not only leverages the pre-trained model’s robust feature extraction capabilities but also significantly reduces computational costs and training time compared to training models from scratch. However, as many pre-trained large vision models are non-public and task-specified datasets may involve highly private samples, the personalized models are a valuable intellectual property for their owners.
II-B Model Stealing Attack
@@
Model stealing refers to unauthorized replication of a victim model’s functionality, which compromises the intellectual property of model owners. Based on the adversary’s access level, existing attacks can be categorized as follows:
@@
White-Box Stealing Attacks. In this attack scenario, the adversary has full access to the model parameters through either public availability (open-source release) or unauthorized internal access. The adversary can directly copy the victim model or fine-tune it to obtain a stolen model.
@@
Black-Box Stealing Attacks. Under this threat model, the adversary can only query the victim model, obtaining the logit outputs. We distinguish two attack scenarios based on the adversary’s data access: (1) Training-Data Assisted Attack: Given access to the training dataset, the adversary can employ knowledge distillation [11] to train a surrogate model that mimics the victim’s input-output behavior. (2) Query-Only Attacks: Without access to any training samples, the adversary can still conduct model stealing via data-free knowledge distillation (DFKD) [44], label-query attack [45, 46, 10], and logits query attack [9, 47, 13]. Specifically, the DFKD method employs an additional generative model to generate data samples for the distillation process. Label-query attacks utilize the victim model to label substitute (unlabeled) samples, which are then used to train their substitute model. Logit-query attacks, on the other hand, obtain a substitute model by minimizing the distance between its predicted logits of the substitute samples and those generated by the victim model.
II-C Defense against Model Stealing Attack
@@
Active Defenses. Active defenses aim to increase the cost or reduce the efficacy of model stealing attacks by increasing the number of queries or reducing the accuracy of the stolen model. Common strategies include introducing perturbations to model outputs or restricting the output form of the victim model. For examples, defenders may apply output probability rounding [9], add noise to logit outputs [48, 49], or only return the most confident label instead of the whole output vector [47]. These methods disrupt the adversary’s ability to accurately mimic the behavior of the victim model, thereby defending against the model stealing attacks. However, these defenses may significantly degrade the performance of the victim models and may even be ineffective against adaptive stealing attacks [17, 25, 50, 24].
@@
Verification-Based Defenses. Generally, model ownership verification aims to verify whether a suspicious model is stolen from the victim model. Existing verification-based methods can be divided into two main categories, including model watermarking and model fingerprinting.
@@
(1) Model Watermarking. Most model watermarking methods conduct verification by modifying the victim model to induce defender-specified misclassification behaviors [14, 15, 16, 17, 18, 19]. For example, backdoor-based verification embeds specific trigger-label pairs into the training data, which lie outside the distribution of the original data. During verification, a model is considered stolen if it misclassifies the triggered samples to the target label, similar to the victim model’s behavior. However, these methods introduce latent vulnerabilities, such as hidden backdoors, which can be exploited maliciously, compromising the model’s functionality and reliability. Additionally, they may be vulnerable to backdoor removal techniques [51, 52, 53], which can erase the embedded watermarks [24].
@@
To the best of our knowledge, MOVE [24] represents the state-of-the-art (SOTA) in harmless model watermarking. Unlike backdoor-based methods, MOVE embeds external features as dataset-specific features into the training set by modifying a subset of training samples using style transfer. Thus, MOVE can identify the stolen model by verifying whether the model has learned the external features. Crucially, these dataset modifications do not alter the labels of the affected data, ensuring the method’s harmlessness. Specifically, considering a $K$-classification problem, the defender pre-defines a style transformer $T:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$, where $\bm{x}_{s}$ is a defender-specified style image. The defender then randomly selects $\gamma\%$ samples ($i.e.$, $\mathcal{D}_{s}$) from the training set $\mathcal{D}$ to generate their transformed version $\mathcal{D}_{t}=\{(\bm{x}^{\prime},y)|\bm{x}^{\prime}=T(\bm{x},\bm{x}_{s}),(\bm{x},y)\in\mathcal{D}_{s}\}$. The victim model $V$ is trained with $\mathcal{D}_{t}\cup\mathcal{D}_{b}$, where $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$. Consequently, the output difference between a suspicious model $S$, $i.e.,$ $S(\bm{x}^{\prime})-S(\bm{x})$, serves as an indicator of potential model theft. However, MOVE is primarily designed for scenarios where the victim model is trained from scratch. As we reveal in Section III, MOVE becomes ineffective for personalized models, as these models struggle to learn external, dataset-specific features during the fine-tuning process since the pre-existing knowledge in foundation models inhibits the learning of external features during fine-tuning.
@@
(2) Model Fingerprinting. Unlike model watermarking methods, model fingerprinting does not require any modifications to the victim model’s training process, ensuring a harmless verification process. Most existing methods [25, 26, 27] verify ownership by leveraging the inherent features learned by the victim model from its original training data. For example, dataset inference [25] is a classical model fingerprinting method that uses the decision boundary of the victim model as a representation of the learned features. Specifically, for each sample $(\bm{x},y)$ in the training set, dataset inference defines the minimum distance to each class ($i.e.$, $\bm{\delta}=(\bm{\delta}_{1},\cdots,\bm{\delta}_{K})$) as the decision boundary of the victim model $V$. The minimum distance $\bm{\delta}_{t}$ to each class $t$ is generated via:
@@
where $d(\cdot)$ is a distance metric ($e.g.$, $\ell^{\infty}$ norm). The defender then selects samples from private and public datasets, using $\bm{\delta}$ as feature embeddings to train a binary meta-classifier $C$, where $C(\bm{\delta})\in[0,1]$ estimates the probability of a sample belonging to the private dataset. To identify stolen models, the defender creates equal-sized sample vectors from private and public samples and conducts a hypothesis test based on the trained $C$. If the confidence scores of private samples are significantly higher than those of public samples, the suspicious model is treated as containing the features from private samples, thereby being identified as stolen model.
Follow-up works [26, 27] aim to improve the characterization of learned inherent features. For example, Peng et al. introduced universal adversarial perturbations (UAPs) [26] to capture the decision boundary of the victim model, which serves as a representation of the learned inherent features. Metafinger [27] leverages meta-training to capture the decision regions of the victim model. Metafinger generates multiple shadow models as meta-data and optimizes specific images through meta-training, ensuring that only models derived from the victim model can correctly classify them. However, model fingerprinting methods are primarily designed for simple model stealing attacks, $e.g.$, direct copying or fine-tuning. These methods exhibit limited robustness against advanced model stealing methods [50, 24], $e.g.$, knowledge distillation [11, 44], making them insufficient for protecting models in complex adversarial scenarios. Additionally, as we reveal in Section III, verification based solely on inherent features may easily lead to misjudgments, especially when the training sets of the suspicious model and the victim model have similar distributions [50, 24]. This is primarily because different models may learn similar common features when their training sets share certain distributional similarities.
@@
III Revisiting Harmless Model Ownership Verification
In this section, we revisit harmless model ownership verification, which includes harmless model watermarking and model fingerprinting methods. We take on two representative approaches as study cases: MOVE [24] for harmless watermarking and dataset inference [25] for model fingerprinting. Specifically, we analyze the limitations of these methods when applied to fine-tuned personalized models.
III-A Preliminaries

### v2

@@ Abstract
Large vision models (LVMs) achieve remarkable performance in various downstream tasks, primarily by personalizing pre-trained models through fine-tuning with private and valuable local data, which makes the personalized model a valuable intellectual property. Similar to the era of traditional DNNs, model stealing attacks also pose significant risks to LVMs. However, this paper reveals that most existing defense methods (developed for traditional DNNs), typically designed for models trained from scratch, either introduce additional security risks, are prone to misjudgment, or are even ineffective for fine-tuned models. To alleviate these problems, this paper proposes a harmless model ownership verification method for personalized LVMs by decoupling similar common features. In general, our method consists of three main stages. In the first stage, we create shadow models that retain common features of the victim model while disrupting dataset-specific features. We represent the dataset-specific features of the victim model by computing the output differences between the shadow and victim models, without altering the victim model or its training process. After that, a meta-classifier is trained to identify stolen models by determining whether suspicious models contain the dataset-specific features of the victim. In the third stage, we conduct model ownership verification by hypothesis test to mitigate randomness and enhance robustness. Extensive experiments on benchmark datasets verify the effectiveness of the proposed method in detecting different types of model stealing simultaneously. Our codes are available at https://github.com/zlh-thu/Holmes.
keywords:
Ownership Verification, Model Fingerprinting, Model Stealing, Deep IP Protection, AI Security
1 Introduction
Pre-trained large vision models (LVMs) are widely used across various applications [VaswaniSPUJGKP17, radford2021learning, gu2024rwkv]. The dominant development paradigm for specific downstream tasks involves personalizing LVMs via fine-tuning on local data. This approach results in a well-performing task-oriented model at a small training cost since the foundation model has already learned a high-quality representation from a large amount of data [wei2023improving, wang2023image, li2025rethinking]. Generally, these personalized models have significant commercial value, and their fine-tuning process involves private and valuable data, making them critical intellectual property for the owners.
@@ 1 Introduction
However, recent studies revealed that the adversary could ‘steal’ (personalized) models even without access to their training samples or model parameters [tramer2016stealing, chandrasekaran2020exploring, carlini2024stealing] via model stealing attacks. For example, the adversary might use the victim model to label an unlabeled dataset or apply knowledge distillation to reproduce the victim model [gou2021knowledge, li2023curriculum, sun2024logit]. Such attacks pose a severe threat to model owners, as these attacks require significantly fewer resources than directly conducting personalization ($e.g.$, fewer computational resources or less training data).
@@
Currently, model ownership verification (MOV) is an important audit strategy to protect the model copyright. It aims to identify whether a suspicious third-party model is stolen from the victim. Existing MOV methods usually fall into two categories: model watermarking and model fingerprinting. Specifically, the model watermarking approaches [ya2023towards, zhu2024reliable, wang2025sleepermark] mostly rely on embedding defender-specified misclassification behaviors, especially via backdoor attacks [li2022backdoor, gao2024backdoor, liang2025vl], into the victim model. However, these methods inevitably introduce potential vulnerabilities, compromising model reliability and limiting practical deployment. Most recently, to address this limitation, MOVE [li2025move] proposed a harmless watermarking scheme that avoids embedding misclassification behaviors. Specifically, MOVE defined harmlessness as the watermarked model should have similar prediction behaviors to the one without any watermark ($i.e.$, benign model trained with the benign dataset). Accordingly, instead of leveraging misclassification behaviors, MOVE modified a subset of training samples using style transfer without manipulating their labels to embed external features as dataset-specific features into the training set. By identifying whether the model has learned the external features, MOVE can provide effective model ownership verification. In particular, since the modifications to the training set do not alter the labels of the affected data, this method remains harmless. However, it is primarily designed for models trained from scratch. As we will reveal in Section 3, MOVE is ineffective for fine-tuned models because the required external features cannot be adequately learned during fine-tuning due to interference from extensive ‘robust features’ learned by the foundation model.
@@
In contrast, model fingerprinting methods [peng2022fingerprinting, yang2022metafinger, shao2025sok] verify ownership by detecting whether the suspicious models retain the inherent features naturally learned by the victim model from its training data. For example, dataset inference [maini2021dataset] uses the decision boundary of the victim model as a representation of its learned features. Unlike watermarking, fingerprinting does not modify the model’s training process, therefore naturally ensuring harmlessness. However, these methods are primarily designed for simple model stealing attacks ($e.g.$, direct copying or fine-tuning), whereas exhibiting limited robustness against advanced stealing methods [maini2021dataset, li2025move]. Besides, as we will show in Section 3, relying solely on inherent features may lead to misjudgment when the victim and suspicious models are trained on similar datasets, as they may learn similar ‘common features’. These limitations raise an intriguing question: Can we design effective and harmless MOV methods for personalized LVMs?
@@
Fortunately, the answer to the above question is positive, although its solution is non-trivial. Motivated by these insights, we aim to reduce misjudgments in model fingerprinting by decoupling common features and isolating dataset-specific characteristics from the model’s inherent representations. We then exploit these decoupled inherent yet dataset-specific features for fingerprinting: the inherent component ensures effective learnability, while the dataset-specific component reduces the likelihood of misjudgment. In contrast to prior fingerprinting approaches that mainly optimize the representation of inherent features [peng2022fingerprinting, yang2022metafinger, shao2025sok], this paper advances a new fingerprinting paradigm that explicitly suppresses the interference of shared common features, which are likely a primary source of misjudgments. In general, our method (dubbed ‘Holmes’) consists of three main stages, including (1) creating shadow models, (2) training ownership meta-classifier, and (3) ownership verification with hypothesis test. Specifically, in the first stage, we create two shadow models, including one poisoned shadow model and one benign shadow model, to decouple the common features and distinguish the learned dataset-specific features of the victim model. In general, the poisoned shadow model retains common features learned by the victim model, while its dataset-specific features are disrupted. As such, we can indirectly capture the dataset-specific features while decoupling the common features via the output difference between these two models. To achieve it, we poison the most well-learned personalization samples (with the lowest loss values) to disrupt dataset-specific features via the label-inconsistent backdoor attack ($e.g.$, BadNets [gu2019badnets]). Intuitively, we choose these samples since they correspond to the model’s propensity to memorize dataset-specific patterns (instead of common features) [ilyas2019adversarial]. We exploit the (label-inconsistent) backdoor attack since it allows a high performance of poisoned models in predicting benign testing samples, thus preserving common features. The benign shadow model is introduced for reference to further mitigate false positives from independent models with related dataset-specific features. In general, this model exhibits dataset-specific features that are similar but distinct from those of the victim model. Accordingly, the output difference between the benign and poisoned shadow models reveals dataset-specific features that are distinct from those of the victim model. To achieve this model, we fine-tune the same foundation model on personalization samples excluding those that were well-learned by the victim model. In the second stage, we train a binary ownership meta-classifier for identifying dataset-specific features of the victim model, based on two output differences: (1) the output differences between the victim model and the poisoned shadow model (for capturing the victim model’s dataset-specific features) and (2) the output differences between the benign shadow model and the poisoned shadow model (for capturing features unrelated to the victim model’s dataset-specific features). In the third stage, we conduct model ownership verification via hypothesis test to mitigate randomness and further enhance robustness. Throughout all stages of our workflow, neither the (released) victim model nor its training process is modified. The proposed method is therefore intrinsically harmless.
@@
The main contribution of this work is four-fold: (1) We revisit the ownership verification methods and reveal the limitations of existing methods for personalized LVMs. (2) We propose a new fingerprinting paradigm that suppresses the interference of shared common features, addressing a primary cause of misjudgments in ownership verification. (3) We develop an effective and intrinsically harmless model fingerprinting method based on feature decoupling, instantiated through the construction of shadow models and hypothesis-test-based verification. (4) We validate the effectiveness of our method on benchmark datasets for simultaneously detecting multiple types of model stealing and further discuss its potential extensions to generative tasks such as image captioning.
2 Related Work
2.1 Large Vision Model
@@
By learning high-quality representations from extensive datasets, large pre-trained vision models (LVMs) have demonstrated remarkable capabilities in comprehending image content and extracting rich semantic information [ShenLTBRCYK22, wang2023image, gu2024rwkv]. These models excel across a wide range of computer vision tasks, including image classification [mauricio2023comparing, gao2024clip, al2025analysing], transfer learning [wei2023improving, wang2023image, gu2024rwkv], object detection [kirillov2023segment, cai2024crowd, wu2025weakly], and semantic segmentation [wang2025reclip++, zhu2025weakclip, xie2025clims++]. Among the most notable LVMs, CLIP (Contrastive Language-Image Pre-Training) [radford2021learning] stands out as a cornerstone in the field. CLIP leverages contrastive learning to align textual and visual information in a shared embedding space, enabling it to perform image classification [mauricio2023comparing, gao2024clip, al2025analysing] and transfer learning [rasheed2023fine] with exceptional accuracy. Pre-trained on a massive dataset of image-text pairs, CLIP is widely adapted as a foundation model for various downstream tasks [mokady2021clipcap, lin2024open, sun2025fairness]. Other prominent LVMs include BEiT (Bidirectional Encoder Representation from Image Transformers) [bao2021beit], which also achieves outstanding performance in various vision tasks [wang2023image].
@@
Currently, the most prevalent application paradigm for LVMs involves personalization on downstream tasks via fine-tuning [wei2023improving, wang2023image, gu2024rwkv]. This process typically begins with a pre-trained foundation model, which has already learned high-quality generalizable features from a large dataset. The foundation model is then adapted to specific tasks by fine-tuning on task-specific datasets. For example, in medical imaging, a pre-trained LVM can be fine-tuned on a dataset of X-ray images to significantly improve disease diagnosis accuracy [lin2024open, sun2025fairness]. Fine-tuning not only leverages the pre-trained model’s robust feature extraction capabilities but also significantly reduces computational costs and training time compared to training models from scratch. However, as many pre-trained LVMs are non-public and task-specified datasets may involve highly private samples, the personalized models are a valuable intellectual property for the owners.
2.2 Model Stealing
@@
Model stealing refers to the unauthorized replication of a victim model’s functionality, thereby compromising the model owner’s intellectual property. Depending on how the adversary interacts with the victim model, existing attacks can be categorized into the following types:
@@
Parameter-Based Stealing. This category involves direct use of the victim model’s parameters, requiring full access to the model through either public release or unauthorized means. The adversary may directly copy the model or fine-tune it to produce a stolen model [maini2021dataset, li2025move].
@@
Distillation-Based Stealing. In this setting, the adversary does not directly access or modify the victim model’s parameters. Instead, the victim model is treated as a teacher model, and a surrogate (student) model is trained to mimic its functionality via knowledge distillation. These attacks are further divided based on the adversary’s data access: (1) Training Data-Assisted Attacks: Given access to the training dataset, the adversary can employ knowledge distillation [li2023curriculum] to train a surrogate model that mimics the victim’s input-output behavior. (2) Query-Only Attacks: Without access to any training samples, the adversary can still conduct model stealing via data-free distillation [yu2023data, liu2024small, tran2024nayer], hard distillation attack [papernot2017practical, jagielski2020high], and soft distillation attack [orekondy2019knockoff, cai2025llava]. Specifically, the data-free distillation method employs an additional generative model to generate data samples for the distillation process. Hard distillation attacks utilize the victim model to label substitute (unlabeled) samples, which are then used to train their substitute model. Soft distillation attacks typically obtain a substitute model by minimizing the distance between its predicted logits of the substitute samples and those generated by the victim model. In particular, some approaches, such as the Distilled Fine-Tuning (DFT) [cai2025llava], extend beyond using logits and exploit intermediate representations (e.g., vision tokens). While this enables more effective knowledge transfer, it comes at the cost of requiring access to the model’s internal features.
2.3 Defense against Model Stealing
@@
Active Defenses. Active defenses aim to increase the cost or reduce the efficacy of model stealing attacks by increasing the number of queries or reducing the accuracy of the stolen model. Common strategies include introducing perturbations to model outputs or restricting the output form of the victim model. For examples, defenders may apply output probability rounding [tramer2016stealing], add noise to logit outputs [lee2018defending, zhang2023apmsa], or only return the most confident label instead of the whole output vector [orekondy2019knockoff]. These methods disrupt the adversary’s ability to accurately mimic the behavior of the victim model, thereby defending against the model stealing attacks. However, these defenses may significantly degrade the performance of the victim models and may even be ineffective against adaptive stealing attacks [jia2021entangled, maini2021dataset, li2025move].
@@
Verification-Based Defenses. In general, model ownership verification aims to verify whether a suspicious model is stolen from the victim model. Existing verification-based methods can be divided into two main categories, including model watermarking and model fingerprinting.
@@
(1) Model Watermarking. Most model watermarking methods conduct verification by modifying the victim model to induce defender-specified misclassification behaviors [jia2021entangled, wang2023free, zhu2024reliable]. For example, backdoor-based verification embeds specific trigger-label pairs into the training data, which lie outside the distribution of the original data. During verification, a model is considered stolen if it misclassifies the triggered samples to the target label, similar to the victim model’s behavior. However, these methods introduce latent vulnerabilities, such as hidden backdoors, which can be exploited maliciously, compromising the model’s functionality and reliability. Additionally, they may be vulnerable to backdoor removal techniques [wang2019neural, chen2024anti, zhu2024neuralsanitizer], which can erase the watermarks [li2025move].
@@
To the best of our knowledge, MOVE [li2025move] represents the state-of-the-art (SOTA) in harmless model watermarking. Unlike backdoor-based methods, MOVE embeds external features as dataset-specific features into the training set by modifying a subset of training samples using style transfer. Thus, MOVE can identify the stolen model by verifying whether the model has learned the external features. Crucially, these dataset modifications do not alter the labels of the affected data, ensuring the method’s harmlessness. Specifically, considering a $K$-classification problem, the defender pre-defines a style transformer $T:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$, where $\bm{x}_{s}$ is a defender-specified style image. The defender then randomly selects $\gamma\%$ samples ($i.e.$, $\mathcal{D}_{s}$) from the training set $\mathcal{D}$ to generate their transformed version $\mathcal{D}_{t}=\{(\bm{x}^{\prime},y)|\bm{x}^{\prime}=T(\bm{x},\bm{x}_{s}),(\bm{x},y)\in\mathcal{D}_{s}\}$. The victim model $V$ is trained with $\mathcal{D}_{t}\cup\mathcal{D}_{b}$, where $\mathcal{D}_{b}\triangleq\mathcal{D}\backslash\mathcal{D}_{s}$. Consequently, the output difference between a suspicious model $S$, $i.e.,$ $S(\bm{x}^{\prime})-S(\bm{x})$, serves as an indicator of potential model theft. However, MOVE is primarily designed for scenarios where the victim model is trained from scratch. As we reveal in Section 3, MOVE becomes ineffective for personalized models, as these models struggle to learn external, dataset-specific features during the fine-tuning process. This is primarily because the pre-existing knowledge in the foundation model inhibits the learning of new external features.
@@
(2) Model Fingerprinting. Unlike model watermarking methods, model fingerprinting does not require any modifications to the victim model’s training process, ensuring a harmless verification process. Most existing methods [maini2021dataset, peng2022fingerprinting, yang2022metafinger] verify ownership by leveraging the inherent features learned by the victim model from its original training data. For example, dataset inference [maini2021dataset] is a classical model fingerprinting method that uses the decision boundary of the victim model as a representation of the learned features. Specifically, for each sample $(\bm{x},y)$ in the training set, dataset inference defines the minimum distance to each class ($i.e.$, $\bm{\delta}=(\bm{\delta}_{1},\cdots,\bm{\delta}_{K})$) as the decision boundary of the victim model $V$. The minimum distance $\bm{\delta}_{t}$ to each class $t$ is generated via:
@@
where $d(\cdot)$ is a distance metric ($e.g.$, $\ell^{\infty}$ norm). The defender then selects samples from private and public datasets, using $\bm{\delta}$ as feature embeddings to train a binary meta-classifier $C$, where $C(\bm{\delta})\in[0,1]$ estimates the probability of a sample belonging to the private dataset. To identify stolen models, the defender creates equal-sized sample vectors from private and public samples and conducts a hypothesis test based on the trained $C$. If the confidence scores of private samples are significantly higher than those of public samples, the suspicious model is treated as containing the features from private samples, thereby being identified as a stolen model.
Follow-up works [peng2022fingerprinting, yang2022metafinger] aim to improve the characterization of learned inherent features. For example, Peng et al. introduced universal adversarial perturbations (UAPs) [peng2022fingerprinting] to capture the decision boundary of the victim model, which serves as a representation of the learned inherent features. Metafinger [yang2022metafinger] leverages meta-training to capture the decision regions of the victim model. Metafinger generates multiple shadow models as meta-data and optimizes specific images through meta-training, ensuring that only models derived from the victim model can correctly classify them. However, model fingerprinting methods are primarily designed for simple model stealing attacks, $e.g.$, direct copying or fine-tuning. These methods exhibit limited robustness against advanced model stealing methods [li2022defending, li2025move], $e.g.$, knowledge distillation [li2023curriculum, tran2024nayer], making them insufficient for protecting models in complex adversarial scenarios. Additionally, as we reveal in Section 3, verification based solely on inherent features may easily lead to misjudgments, especially when the training sets of the suspicious model and the victim model have similar distributions [li2022defending, li2025move]. This is primarily because different models may learn similar common features when their training sets share certain distributional similarities.
@@
3 Revisiting Harmless Model Ownership Verification
This section revisits harmless model ownership verification in the context of fine-tuned personalized models, covering both harmless model watermarking and model fingerprinting. We take on two representative approaches as study cases: MOVE [li2025move] for harmless watermarking and dataset inference [maini2021dataset] for model fingerprinting. Our analysis specifically focuses on their limitations when applied to fine-tuned personalized models.
3.1 Preliminaries